OpenStudy (anonymous):

Suppose you buy a CD for $750 that earns 4% APR and is compounded quarterly. The CD matures in 3 years. Assume that if funds are withdrawn before the CD matures, the early withdrawal fee is 3 months' interest. What is the early withdrawal fee on this account?

OpenStudy (tkhunny):

This is a somewhat ambiguous question. "3 months' interest" is NOT well-defined. In 3 years, there are 12 obvious 3-month periods, 34 three-month periods if we insist on whole months, and infinitely many 3-month periods if we completely liberalize the definition. i = 0.04 j = i/4 = 0.01 r = 1+j = 1.01 Anyway, I think it is most likely possibilities are: 1) The early withdrawal penalty is intended to be the most punitive. This would have to be the interest paid in the LAST three months of the original term. \(750r^{12}-750r^{9}\) 2) The early withdrawal penalty is intended to be easiest to calculate and easiest to justify. This would have to be the interest paid in the FIRST three months of the original term. \(750r-750 = 750j\)

OpenStudy (anonymous):

APEX: $7.50